OpenStudy (anonymous):

can someone explain how I do this? please! Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)).

OpenStudy (anonymous):

Start by straight out replacing the g(x) so you get f(10x - 6)

OpenStudy (anonymous):

ok now what?

OpenStudy (anonymous):

This means that for f(x) = -4x + 7 We replace the x next to the -4 with (10x - 6)

OpenStudy (anonymous):

You get... using the proper notation f(g(x))= f(10x-6) = -4(10x-6) +7

OpenStudy (anonymous):

whats the first step in solving this? step by step PLease

OpenStudy (anonymous):

The first thing you do is replace the g(x), to get f(10x-6)

OpenStudy (anonymous):

then

OpenStudy (anonymous):

Now using what we know about plugging values into functions... we look at f(x) and we get f(10x-6) = -4(10x-6) + 7

OpenStudy (anonymous):

All that is left is to simplify

OpenStudy (anonymous):

ok ill try that thanks

OpenStudy (anonymous):

Let me know if you need some help with that

OpenStudy (anonymous):

so I simplified the right side and got f(10x-6) = -40+30 ??

OpenStudy (anonymous):

-40x+30**

OpenStudy (anonymous):

Hold on was afk

OpenStudy (anonymous):

Pretty close Just your algebra It is -40x + 31